What is the formula for c speed of light in terms of free space coefficient

Grade:6

1 Answers

Mahima Kanawat
5 years ago
Dear student 
C = 1/√permittivity of free space *vacuum permeability.
C =  (ε'×μ') ^-1/2 or c =1/ (ε'×μ') 1/2
WITH REGARDS 
MAHIMA 
ASKIITIANS FORUM EXPERT